Lab Experiment: Determining the Calorimeter Constant

NEED HELP WITH THE FOLLOWING BOLD TEXTS, PLEASE ANSWER ALL AND NOT PARTIAL. THANK YOU

Trial 1                  Trial 2

A. Volume of room temperature water, mL                                                     75mL                   100mL

B. Volume of hot water, mL                                                                             75mL                   100mL

C.Initial temperature of room temperature water, °C   23.3°C 24.5°C

D. Initial temperature of hot water, °C                                                             20.2°C 22.2°C

E. Final (max.) temperature of mixture, °C                                                     19.5°C 21.8°C

F. ΔT of room temperature water, °C           ________           ________

G. Mass of room temperature water, g    ________           ________

(assume a density of 1.0 g/mL)

H. qrt, J ________           ________

In the space below, show your work for calculating qrt for trial 1

I. ΔT of hot water, °C                                                                                       _______             ________

J. Mass of hot water, g _______             ________

K. qhot, J    ________           ________

In the space below, show your work for calculatimg qrt for trial 1

L. Heat absorbed by calorimeter, J     ________           ________

M. Ccal, J/°C                                                                                                        ________           ________

In the space below, show your work for calculating Ccal for trial 1.

N. Average Ccal                                                                                                                 ________
Answer #1
Trial 1 Trial 2
A Volume of room temperature water, mL 75 100
B Volume of hot water, mL 75 100
C Initial temperature of hot water, ^\circ C 23.3 24.5
D

Initial temperature of room temp water, ^\circ C

20.2 22.2
E final temeprature of mixture, ^\circ C 19.5 21.8
F \Delta T of hot water, ^\circ C 3.8 2.7
G Mass of hot water, g 75 100
H hot, heat lost by hot water, J 1193.01 1130.22
I \Delta T of room temp water, ^\circ C 0.7 0.4
J Mass of room temp water, g 75 100
K q_{rt}, heat gained by room temp water, J 209.765 167.44
L Hear absorbed by calorimeter, J 983.245 962.78
M CCal, J/C 258.74 356.58
N Average Ccal, J/ C 307.66

note: In your data, temp of hot water is lower than room temp water, which is certainly a mistake. Hence I have duly exchanged the data for room temp water and hot water.

F.

Change in temperature hot water,

For Trial 1

\Delta T_1 = |T(after \ mixing) - T(room \ temp \ water)| = |19.5 - 23.3| =3.8 ^\circ C

For Trial 2

\Delta T_2 = |T(after \ mixing) - T(room \ temp \ water)| =| 21.8 - 24.5| =2.7 ^\circ C

G.

Mass of hot water.

For trial 1

Volume of room temp water = 75 mL

Density = 1g/mL

hence, mass = Volume* Density = 75 mL * 1 g/ mL = 75 g

For Trial 2

Volume of room temp water = 100 mL

Density = 1g/mL

hence, mass = Volume* Density = 100 mL * 1 g/ mL = 100 g

H.

Heat lost by hot water.

For trial 1

Heat gained by the room temp water is given by the formula

q_{rt}= m \times C \times \Delta T = 75 \ g \times 4.186 \ J \ g^{-1} \ ^\circ C^{-1} \times 3.8 ^\circ C = 1193.01 \ J

For Trial 2

q_{rt}= m \times C \times \Delta T = 100 \ g \times 4.186 \ J \ g^{-1} \ ^\circ C^{-1} \times 2.7 ^\circ C = 1130.22 \ J

I.

\Delta T for room temp water

For Trial 1

\Delta T_1 = |T(after \ mixing) - T(hot \ water)| = |19.5 - 20.2| =0.7 ^\circ C

For Trial 2

\Delta T_1 = |T(after \ mixing) - T(hot \ water)| = |21.8 - 22.2| =0.4 ^\circ C

J.

Mass of room temp water.

For trial 1

Volume of room temp water = 75 mL

Density = 1g/mL

hence, mass = Volume* Density = 75 mL * 1 g/ mL = 75 g

For Trial 2

Volume of room temp water = 100 mL

Density = 1g/mL

hence, mass = Volume* Density = 100 mL * 1 g/ mL = 100 g

K.

For Trial 1

Heat gained by room temp water is given by

q_{hot}= m \times C \times \Delta T = 75 \ g \times 4.186 \ J \ g^{-1} \ ^\circ C^{-1} \times 0.7 ^\circ C = 209.765 \ J

For Trial 2

q_{hot}= m \times C \times \Delta T = 100 \ g \times 4.186 \ J \ g^{-1} \ ^\circ C^{-1} \times 0.4 ^\circ C = 167.44\ J

L.

The heat lost by the hot water is gained by both the room temp water and calorimeter.

hence,

q_{cal} + q_{rt} = q_{hot} \\ \Rightarrow q_{cal} = q_{hot} - q_{rt}

For Trial 1

q_{cal}= 1193.01 \ J - 209.765 \ J = 983.245 \ J

For Trial 2

q_{cal}= 1130.22 \ J - 167.44 \ J = 962.78 \ J

M.

The temperature change for the calorimeter is same as that of the room temp water.

Also,

q_{cal}= C_{cal} \times \Delta T \\ \Rightarrow C_{cal} = \frac{q_{cal}}{\Delta T}

For Trial 1

\Delta T = 3.8 ^\circ C

C_{cal} =\frac{983.245 \ J}{3.8 \ ^\circ C} = 258.74 \ J/ ^\circ C

For Trial 2

\Delta T = 2.7 ^\circ C

C_{cal} =\frac{962.78 \ J}{2.7 \ ^\circ C} =356.58 \ J/ ^\circ C

N

Average C_cal =

\frac{258.74+356.58}{2} = 307.66 \ J/ ^\circ C